# Split Gibbs Discrete Diffusion Posterior Sampling

### v1

@@ 4.1 Experimental Setup
SG-DPS $\textbf{20.17}_{\pm 3.47}$ 91.2 $\textbf{17.25}_{\pm 3.82}$ 79.4
Table 2: Quantitative results for XOR and AND problems on discretized MNIST. We report the mean and standard deviation of PSNR and class accuracy across 1k generated samples. The class accuracy is computed with a surrogate classifier trained on the training set. SG-DPS demonstrates superior performance on both tasks.
4.2 Controlled Study using Synthetic Data
We first evaluate the accuracy of SG-DPS in posterior sampling on a synthetic task where the true posterior distribution is accessible. Let ${\mathbf{x}}\in\mathcal{X}^{D}$ follow a prior distribution $p({\mathbf{x}})$, obtained by discretizing a Gaussian distribution $\mathcal{N}(0,\sigma^{2}{\bm{I}}_{D})$ over an $D$-dimensional grid. We define a forward model by ${\bm{G}}({\mathbf{x}})=\|f({\mathbf{x}})\|_{1}$ where $f({\mathbf{x}})$ maps ${\mathbf{x}}$ to $\mathbb{R}^{D}$, and likelihood function is defined as $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}|/\sigma_{\mathbf{y}}).$
Figure 3: Sampling results of the XOR task on the discretized MNIST dataset.  SG-DPS faithfully recovers the structural information of the ground truth signal.
Figure 4: Diversified samples when the measurement ${\mathbf{y}}$ is sparse. We present 10 samples generated by SG-DPS when solving an MNIST inpainting task with sparse information. The class histogram is counted on 100 samples, showing SG-DPS generating digits of $1,4,7$ and $9$ that are consistent with the measurement. The ground truth distribution is calculated by counting all test samples with measurement error less than a threshold.
We choose $\sigma_{\mathbf{y}}=1$ specifically for synthetic data experiments. We draw 10k independent samples with each algorithm and compute the frequency map for the first two dimensions of ${\mathbf{x}}$. We measure the Hellinger distance and total variance distance between the empirical distribution and ground truth posterior distribution. As shown in Table 1, our method achieves substantially more accurate posterior sampling compared to baseline methods (by nearly a factor of 2x over SVDD-PM).
We visualize the generated samples in Figure 2 by projecting them onto the first two dimensions and computing their heatmaps. When the number of dimensions is small, e.g., $D=2$, all methods approximate the true posterior distribution well. However, as the dimensionality increases, DPS and SVDD-PM struggle to capture the posterior and tend to degenerate to the prior distribution. While SG-DPS consistently outperforms baseline methods in terms of Hellinger distance, it also shows a slight deviation from the posterior distribution as the dimensionality increases.
4.3 Discrete Image Inverse Problems
We evaluate on a discretized image domain. Specifically, we convert the MNIST dataset (LeCun et al., 1998) to binary strings by discretizing and flattening the images. We then train a discrete diffusion prior on 60k training data.
@@ 4.3 Discrete Image Inverse Problems
As examples of linear and nonlinear forward models on binary strings, we consider AND and XOR operators. We randomly pick $\gamma D$ pairs of positions $(i_{p},j_{p})$ over $\{1,\dots,D\}$, and compute:
${\bm{G}}_{\mathrm{AND}}({\mathbf{x}})=[{\mathbf{x}}_{i_{p}}\wedge{\mathbf{x}}_{j_{p}}]_{p=1,\dots,\gamma D},$
${\bm{G}}_{\mathrm{XOR}}({\mathbf{x}})=[{\mathbf{x}}_{i_{p}}\oplus{\mathbf{x}}_{j_{p}}]_{p=1,\dots,\gamma D}.$ (22)
@@
These logical operators can be challenging to approximate using surrogate functions to guide diffusion sampling.
@@
We assume $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|_{0}/\sigma_{\mathbf{y}})$ for all experiments. We use 1000 binary images from the test set of MNIST and calculate the peak signal-to-noise ratio (PSNR) of the reconstructed image as a metric. Furthermore, we train a simple convolution neural network on MNIST as a surrogate and report the classifier accuracy of the generated samples. As shown in Table 2, SG-DPS outperforms baseline methods by a large margin in both XOR and AND tasks. We present the samples generated by SG-DPS for the XOR task in Figure 3. The reconstructed samples are visually consistent with the underlying ground truth signal, which explains the high class accuracy of our method shown in Table 2.
@@
Sample diversity. Furthermore, we demonstrate that SG-DPS generates diversified samples when the measurement ${\mathbf{y}}$ is sparse. For example, when a digit is masked with a large box, as shown in Figure 4, the measurement lacks sufficient information to fully recover the original digit. In this scenario, SG-DPS generate samples from multiple plausible modes, including digits 1,4,7 and 9. This highlights the ability of our method to produce diverse samples from the posterior distribution while preserving consistency with the measurement.
@@ 4.4 Monophonic Music
We conduct experiments on monophonic songs from the Lakh pianoroll dataset (Dong et al., 2018). Following the preprocessing steps in (Campbell et al., 2022), we obtain monophonic musical sequences of length $D=256$ and vocabulary size $N=129$. The ordering of musical notes is scrambled to destroy any ordinal structure.

### v2

@@ 4.4 Discrete Image Inverse Problems
SGDD $\textbf{20.17}_{\pm 3.47}$ 91.2 $\textbf{17.25}_{\pm 3.82}$ 79.4
@@
Setup. We evaluate on a discretized image domain. Specifically, we convert the MNIST dataset [25] to binary strings by discretizing the images, and train a discrete diffusion prior on 60k training data. As examples of linear and nonlinear forward models on binary strings, we consider AND and XOR operators. We randomly pick $\gamma D$ pairs of positions $(i_{p},j_{p})$ over $\{1,\dots,D\}$, and compute:
${\bm{G}}_{\mathrm{AND}}({\mathbf{x}})=[{\mathbf{x}}_{i_{p}}\wedge{\mathbf{x}}_{j_{p}}]_{p=1,\dots,\gamma D},\ {\bm{G}}_{\mathrm{XOR}}({\mathbf{x}})=[{\mathbf{x}}_{i_{p}}\oplus{\mathbf{x}}_{j_{p}}]_{p=1,\dots,\gamma D}.$ (15)
@@
These logical operators are challenging to approximate using surrogate functions to guide sampling.
@@
Evaluation metrics. We use 1k binary images from the test set of MNIST and calculate the peak signal-to-noise ratio (PSNR) of the reconstructed image. Furthermore, we train a simple convolutional neural network on MNIST as a surrogate and report the classifier accuracy of the generated samples.
Results. As shown in Table 4, SGDD outperforms baseline methods by a large margin in both XOR and AND tasks. We present the samples generated by SGDD for the XOR task in Section D.2. The reconstructed samples are visually consistent with the underlying ground truth signal, which explains the high class accuracy of our method shown in Table 4.
@@
Sample diversity. Furthermore, we demonstrate that SGDD generates diversified samples when the measurement ${\mathbf{y}}$ is sparse. For example, when a digit is masked with a large box, as shown in Fig. 3, the measurement lacks sufficient information to fully recover the original digit. In this scenario, SGDD generate samples from multiple plausible modes, including digits 1,4,7 and 9. This highlights the ability of our method to produce diverse samples from the posterior distribution while preserving consistency with the measurement.
Higher dimensional images. To demonstrate the scalability of SGDD, we further evaluate SGDD on a higher-dimensional dataset, FFHQ 256 [21] in Section D.1, where we solve the super-resolution task by operating in the discrete latent space of a pretrained vector-quantized diffusion model.
@@ 4.5 Monophonic Music
Table 5: Quantitative evaluation of infilling monophonic music sequences.
$\gamma=40\%$ $\gamma=60\%$
Hellinger $\downarrow$ Outliers $\downarrow$ Hellinger $\downarrow$ Outliers $\downarrow$
SVDD-PM $0.465_{\pm 0.129}$ $0.112_{\pm 0.098}$ $0.529_{\pm 0.137}$ $0.114_{\pm 0.119}$
SMC $0.492_{\pm 0.143}$ $0.152_{{\pm}0.138}$ $0.552_{\pm 0.131}$ $0.164_{\pm 0.140}$
SGDD $\textbf{0.126}_{\pm 0.098}$ $\textbf{0.003}_{\pm 0.002}$ $\textbf{0.244}_{\pm 0.164}$ $\textbf{0.060}_{\pm 0.138}$